Category Archives: LCR Meters

HP 4192A LF Impedance Meter: Isolated power supply fixed

Working through the guts of the 4192A, I tried to operate the A8 assembly, which has several +-15 V isolated power supplies (for floating voltages, etc.), and a +-40 V bias supply, set by a pretty cerdip DAC.

Basically, the assembly consists of a set of switched power converters, converting the +-15 V input, to isolated +-15 V output. Running at about 900 kHz to 1 MHz.

The assembly is double shielded, to interference with the impedance measurements. It has a set of coils and capacitors, a number of parts! Definitely not a cheap unit to build at the time.

The issue – when connecting it to the 4192a power supply, it loads down the +-15 V rails, and the power supply assembly A7 of the 4192a soon goes into shut-down. Not a good sign.
Analyzed the circuit by touch, almost burning my finger – the two main switching transistors of the 4th supply are glowing hot!

Did some measurements, and the transistors, as well as two capacitors are dead. The other capacitors, I removed them because you never know, and better to fix it now than later.

Interestingly, the capacitors HP selected were 85°C rated, I decided on some Panasonic EB series, 105°C rated, low ESR-long lifetime caps. Hope these will last another 30 years of occasional service.

The transistors – difficult to find good substitutes. I believe some other reasonably high current NPN fast transistors like SS9050 or similar would work, but I don’t want to take chances and found some new-old-stock 2N3725 on eBay. Only it will take a while for these to arrive in Japan!
For the time being, I desoldered the working transistors of the #3 supply, to get the #4 supply going for the tests.

Surprisingly, one of the transformers has a damaged pot core, K5 ferrite, difficult to get. Strangely, there were no loose pieces inside the shields, and no traces of earlier repairs. Can it be that HP fitted a slightly defective transformer? But as the unit is working and outputting plenty of power, the core should be good enough.

To fix it, or at least make sure that it is not rattling or desintegrating any further, we have a choice of epoxy glue, super glue, or silicone. I decided on the latter, a 704 type head conductive silicone. It is available really, cheap, and it is good stuff, because it can be removed later without destroying the coil, should I ever come across some K5 pot cores.

All fixed and stable.

A quick test – powering from a current limited lab supply (about 80 mA on 5 V, about 0.11 Amps on +15 and -15 V with no load are needed for the A8) to avoid destruction.

Fortunately, the A8 assy is working again. I didn’t test the DAC, but no reason to believe it would not work.

Now just waiting on on the 2n3725A transistors, and then, it can be all put togethers with many screws and pieces of shielding metal.

HP 4192A LF Impedance Analyzer: some issues fixed, more unveiled

After a quick trip to Germany (and with a set of spare parts in my suitcase), I directed my attention to the sick 4192A again.

(1) First, the A7 power supply assembly. Fixed the defective signal diode, and ceramic capacitor. Discovered that the line frequency signal is not working, and that the 5 V supply (analog section) linear regulator isn’t regulating, but passing through about 6 volts. Also noticed that the digital 5 V supply is not working (running at 2-3 volts).

(2) After about 20 seconds of operation, there is some smell around the A7 power supply assembly. The input caps (10 uF, 350 volts) running hot.

(3) No display at all. Is the controller assy dead? Or the EPROMS corrupted?

Let’s tackle it step by step.

Fortunately, we are not alone here, the same instrument had similar issues elsewhere – seems the NiCd batteries weren’t fit for the purpose, and have all leakage issues, as can be seen in below screenshot of a Japanese blogger.

Unsoldered the 10 uF 350 V caps of the A7 power supply – all the positive connections are leaking. Fortunately, not as corrosive as the NiCd electrolyte.

Strangely, HP fitted 85 C capacitors, rather than long life types.

On the picture you can see the temporary fix with some capacitors I had handy. And, you can see the Y-rated capacitors (15 n) RIFA cracked capacitors replaced by WIMA brand new parts.

For replacement of the 10 uF Chemi-Con we will use CFX series capacitors.

These are rated for quite severe ripple current, which is necessary because that’s their purpose in the A7 circuit.

Of course, 2 parallel capacitor 10 uF will have better ESR than a single 22 uF, but fair enough for test purposes.

The digital supply, surprisingly, has no regulation, neither by linear nor by switchmode action (the switchmode circuit is regulated by an independent sense winding on the transformer (running at 29.5 kHz).

The digital ground is connected to analog ground by a 10 Ohm resistor for noise isolation. Other than that, the circuit is rather simple and quickly found the defective part – surprisingly, one of the Schottky rectifiers, a 20FQ040 diode. Maybe it overheated, or it failed just because of age.

For its time, 1980s, it is a respectable diode, with very low forward voltage.

Unfortunately, replacements in DO-4 bolt-mount style Schottkys are very expensive and hard to get. Maybe there are some back home in Germany, in the junk pile waiting to be desoldered, but not sure about it, and no such diodes here in Japan. Anyway, after some measurement, the current of the digital supply will be somewhere around 2.5 to 3 Amps, not too much, for any common switchmode supply rectified diode. Decided on a SBL3040PT. 3 Amps at 0.2 Volts will be less than 1 Watt of dissipation, say, about 40 K temperature rise of a SBL3040PT double diode (such packages have about 40 K/W junction to ambient thermal resistance, probably I will fit a small metal part as heatsink (the existing heatsink may be used after drilling a mounting hole…).

Another issue fixed, a corroded spacer for the 5 V analog supply voltage-limiting Zener (a power Zener, still good!). The spaced could have been cleaned, but not worth the hazzle, so I machined a new one from aluminum alloy.

To the mains sensing and synchronizing circuit – some probing with a scope showed that the U1 comparator (a HP numbered LM339) didn’t work. Replacing it was no easy task, because this area had been affected by the NiCd electrolyte, rendering the solder pads difficult to desolder. Scratching off the corrosion layer from the solder with a screwdriver helped to get some contact with the head, and to finally desolder.

Further probing also showed a few damages to resistors, the wires had come off the main body (seems the corrosion damaged the weld between the resistor case, and the wire).

A quick test – there is a clean sync signal! About 120 Hz, double the mains frequency in this part of Japan. Perfect.

With these basics fixed (except the digital 5 V supply – still waiting for delivery of the diode – just fitted a temporary diode for test purposes), let’s have a look at the CPU board, which is a marvelous piece of engineering, with may TTL, EPROMs, RAMs, etc. – let’s hope we don’t need to fix this complex assembly.

Checking the clock and reset lines – no clock present! After some study of the manual, I understood that the CPU clock is generated from a 20 MHz signal, from the 40 MHz VCO and reference frequency assy. Why is there no 20 MHz? Easy answer, the referency assy has no power except -15 V.

Reason: corroded Molex connector, and broken contacts within. Not easily seen from the outside, but clearly, there is no contact.

Fixed the contacts, at least a few (and ordered more!), and voila, the 20 MHz and 40 MHz are back. No issues with any of the main counters (e.g., keyboard and display scan), good activity on the address bus, etc.

For proper tests, removed the CPU/controller assembly A6 from the unit, and powering it with a lab power supply, to make sure it gets stable power.

After all this, quite some relieve! The unit is starting up, at least as much as it can be told at this point (all the analog circuits disconnected from the A6 controller board – need to first fix their power connections). The startup passes all the RAM and EPROM tests, great! And finally stops at E-50, which means, it can’t find the line sync frequency – which is no surprise, because this signal is currently disconnected.

Next steps will be – (1) Fix the Molex connectors. Quite difficult because these are crimp connections, and the wired have some corrosion making them difficult to solder. (2) Finalize the A7 repairs – the 10 uF capacitors, the Schottky for the digital supply, test the digital supply. (3) Then proceed to the start up and repair-adjustment of the analog circuits, and synthesizer section, many, many, complex circuits, but these don’t have any visible corrosion or damage. Fingers crossed.

HP 4276A LCR Meter: several issues, and several animals found

The HP Yokogawa (Made in Japan) LCR meters are still very useful, these are 0.1% basic accuracy LCR meters – with various test frequencies from 100 Hz to 20 kHz. And, you can apply DC bias as needed to test ceramic capacitors, diodes, etc.

I would this meter for a marginal price, in Japan. It arrived in good overall state and well packaged, but very dirty inside. A quick review:

(1) Power supply seems OK, but the fan is not working.
(2) Display shows overrange and self test shows errors in the analog circuit. In some ranges, I get a reading but it is far off.
(3) Dirt, dirt, dirt – this instrument needs to be fully disassembled and cleaned, no doubt, and the front panel needs special treatment to remove all stickers and residues.

The top view – a lot of dust, and of the sticky kind.

In the corners and gaps, a wasp, and a dried frog. Probably, this instrument has been stored in some shed?

The main board – first, brushed of the dirt, then soaked it in 50% isopropanol. Take care of the fan hybrid, which is a thin ceramic plate – overall, a rather delicate construction. Not sure why they did’t just use regular components rather than a hybrid for the fan drive – there is a huge amount of unused space in the case!

The fan, it is a brushless fan, Buehler, which needs an external driver to commutate the phase.

The test setup – with an external 8 Volts supply. Turns out, the fan starts if we move it a bit. Probably, some issue with the driver.

A good phase – you clearly see the switching action of the transistor, grounding the phase wire for some time to move the fan forward.

After some probing, it is clear that one of the 3 phases isn’t working!

Tested all the timing capacitors of the driver – all good!

But – one of the transistors, SMD, marked “B15” is not OK! B15, this is a 2n2222a equivalent transistor – watch out, there are different pinouts.

Unfortunately, I had no SMD NPN transistor of suitable kind at hand, so I soldered in a 2N5551, for test purposes.

It is working – the transistor has a bit less strong/fast switching action, but no problem, the fan is working again!

Next – we have to fix the analog circuit. This is a nice board, with several hybrids. First, we check the power supplies – HP designed this to have all individual +-12 Volt regulators, for each sub-section (all sections are nicely labeled, and clearly identified).

After some difficult probing – found one bad 12 V regulator! It even had some traces of corrosion – maybe it was running hot or some other failure.

Some test – still not working – at least not in all ranges. Seems the input amplifier is clipping the signal. Clearly, something is wrong with the range switching.

… about 1 hour later, traced the signal back to the range selector logic, one bit is not getting through – it got stuck at the edge connector of the CPU board – the edge connector shows some residue and oxidation. Cleaned it thoroughly with an eraser, and alcohol.

Finally, self test passed!

… testing a 22 nF capacitor.

It is a really great instrument to test ceramic capacitors, and the impact of DC bias on their capacity!

As a last measure, inspected all the X and Y rated capacitors – it good practice to do so with any old instrument. These are safety capacitors, and they often fail with age and use.

The power supply cover is connected to the mains with a Y rated capacitor, and these don’t look good.

.. fitted new Y capacitors – feeling much safer now!